I am utterly confused how to get 24/7. The closest that I got is 24. I do not understand how to get this answer. Can someone make a proof to help me understand?
In the system of equations
x+2y = 5
4x-6y = 9
(a) x = 0
(b) x = 11/4
(c) x = 2
(d) x = 24/7 <--- This is the answer. How do you get it?
(e) x= 5

Answer by jim_thompson5910(35256)

Start with this system

x+2y = 5
4x-6y = 9


Multiply both sides of equation 1 by 3


3*(x+2y) = 3*5
4x-6y = 9


Multiply and distribute to get this:


3x+6y = 15
4x-6y = 9


Now add the equations


3x+6y = 15
4x-6y = 9
-----------
7x+0y = 24


So we have...


7x = 24

x = 24/7

------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------